Compute the following to get the equilibrium price and quantity.


1) If Qd = 30-2P and Qd is 4. What is now P?


2) If the Demand Function is Qd = 30-2P and P is 2. What is now Qd?


3) Qd= 30-5P


4) Qs= 0+5P



1
Expert's answer
2022-01-18T13:01:41-0500

1) If Qd = 30 - 2P and Qd is 4, then:

30 - 2P = 4,

P = 13.

2) If the Demand Function is Qd = 30 - 2P and P is 2, then Qd = 30 - 2×2 = 26 units.

3-4) Qd = Qs,

30 - 5P = 0 + 5P,

10P = 30,

P = 3.